# Study of the decay pattern of $f_{0}(1370)$ as a $\kappa\bar{\kappa}$ molecular state

### v1

@@ II Formalism
The constant $1/2$ are the symmetry factor on account of the two identical particles $\pi^{0}$.
For the $\sigma\sigma\to 4\pi^{0}$ process, the amplitudes $\mathcal{M}^{S}_{\sigma}$ and $\mathcal{M}^{V}_{\sigma}$ for $2\pi^{+}2\pi^{-}$ could be directly adopted but need to further considering all possible exchanges of the $\pi^{0}$ as mentioned before.
The the decay width of the $4\pi^{0}$ state is
@@
The constant $1/4!$ are the symmetry factor on account of the four identical particles $\pi^{0}$.
For the $K\bar{K}\pi\pi$ channel from the tree-level decay $f_{0}(1370)\to\kappa\bar{\kappa}\to(K\pi)(\bar{K}\pi)$, as shown in Fig. 1(d), there are six different final states specifically. The $K^{+}K^{-}\pi^{+}\pi^{-}$ and $K^{0}\bar{K}^{0}\pi^{0}\pi^{0}$ can only come from the $\kappa^{0}\bar{\kappa}^{0}$ component; the $K^{0}\bar{K}^{0}\pi^{+}\pi^{-}$ and $K^{+}K^{-}\pi^{0}\pi^{0}$ can only come from the $\kappa^{+}\kappa^{-}$ component; while the $K^{0}K^{-}\pi^{+}\pi^{0}$ and $K^{+}\bar{K}^{0}\pi^{0}\pi^{-}$ can come from both the $\kappa^{0}\bar{\kappa}^{0}$ and $\kappa^{+}\kappa^{-}$ components.
@@
where $s_{ab}$ and $s_{cd}$ are respectively the invariant mass squares of the $K\pi$ pairs coming from $\kappa$ and $\bar{\kappa}$. The amplitudes of other final states can be obtained by replacing the couplings in $\mathcal{M}_{K^{+}K^{-}\pi^{+}\pi^{-}}$ and carefully including all the possible diagrams. Note that, the combinations of $K\pi$ are definite for all the $K\bar{K}\pi\pi$ channels except for $K^{+}K^{-}\pi^{0}\pi^{0}$ and $K^{0}\bar{K}^{0}\pi^{0}\pi^{0}$ due to the two identical particles $\pi^{0}$. Hence, for these two final states, we need to include another diagram obtained by exchanging these two $\pi^{0}$ and when calculating the decay width, the symmetry factor $1/2$ should be considered on account of the double counting in the phase space.
@@ III Numerical results and discussions
Firstly, we present the results calculated with the initial coupling $g_{F\kappa\bar{\kappa}}$ obtained by Weinberg compositeness. The partial decay widths of different channels including $K\bar{K}$, $\pi\pi$, $\eta\eta$, $4\pi$ and $K\bar{K}\pi\pi$ are shown in Table 2. These are the primary strong decay channels of $f_{0}(1370)$, which are approximately summed up to the total width of $f_{0}(1370)$: the $\Gamma_{F}$ in Table 2. We present results for two different couplings $g_{\kappa K\eta}$ and $g^{\prime}_{\kappa K\eta}$. For clarity, we denote the result associated with $g^{\prime}_{\kappa K\eta}$ as $\Gamma^{\prime}$. Note that we sum up the $\sigma\sigma$ and $\rho\rho$ contributions to be the total $4\pi$ width, assuming that the $4\pi$ events from other intermediate states are negligible. In our framework, the $f_{0}(1370)$ decaying to the $K\bar{K},\pi\pi,\eta\eta$ and $4\pi$ channels are via the triangle loop diagrams, which are dependent on the cut-off parameters $\alpha$. Hence, we also present the results with different $\alpha$ in Table 2, to test the cut-off dependence. From Table 2, we can see that when $\alpha$ varies from $2$ to $3$, the decay widths are relatively stable. These decay channels have been observed in experiments except the $K\bar{K}\pi\pi$ channel, which lacks measurements up to now.
Regardless of whether the components are broad or narrow, they are nearly free particles inside the molecular state in a weakly binding situation. Decay through the almost free components implies that the width of the bound state is expected to be of the same order of magnitude as the width of its components. This argument is consistent with the case of $T^{+}_{cc}$, whose width is almost saturated by the decay via $D^{*}D\to DD\pi$ and is analyzed to be about $48$ keV [67], and the width of $D^{*}$ is $(83.4\pm 1.8)$ keV [8]. Hence, aligning with this argument the broad state $f_{0}(1370)$ could be reasonably suspected to be a molecular state composed of $\kappa\bar{\kappa}$. The pole width $\Gamma=-2\,\mathrm{Im}\,\sqrt{s_{\text{pole}}}$ for $f_{0}(1370)$ is estimated to be around $500$ MeV according to its pole position, and its Breit-Wigner width is measured to be between $200$ MeV and $500$ MeV, depending on the models and the reactions. However, as shown in Table 2, with the $g_{F\kappa\bar{\kappa}}=13.4$ GeV determined by the Weinberg criterion and the cut off parameter $\alpha=2$, the sum of all decay channels results in a total width $\Gamma_{F}=39.6/\Gamma^{\prime}_{F}=37.7$ MeV, which is much smaller than all the measured widths of $f_{0}(1370)$ collected in the RPP [8]. Recall that the Weinberg criterion is derived from the case of weakly bound state consisting of two stable or nearly stable components. In contrast, in our case the component $\kappa$ is broad as about $400$ MeV, which implies that the initial decay coupling $g_{F\kappa\bar{\kappa}}$ might still be underestimated even though a complex mass are used in the Weinberg criterion.
@@
We realize that it is complex to figure out the coupling between $f_{0}(1370)$ and its broad components, let alone this molecular state might essentially be composed of four stable particles $K\bar{K}\pi\pi$. Therefore, we take this initial couplings as a parameter to fit the total width of $f_{0}(1370)$ measured in experiments. In Fig. 2, we show the total width of $f_{0}(1370)$ at $\sqrt{s}=1.37$ GeV with $g_{F\kappa\bar{\kappa}}$ as a parameter. We also present the results with different cut-off parameters $\alpha$ from $1\sim 3$. Notice that this initial couplings is a common coupling for all the decay channels, its change will not affect the ratios among the decay channels. In other words, the decay pattern is independent of this initial coupling. The light purple band in Fig. 2 indicates the range of the Breit-Wigner width of $f_{0}(1370)$ measured in experiments. The vertical red dashed line indicates the value of $g_{F\kappa\bar{\kappa}}$ from the Weinberg criterion using a complex threshold. We can see that when cut-off parameter $\alpha=1$, even the coupling $g_{F\kappa\bar{\kappa}}$ is taken as large as $45$ GeV, the total width $\Gamma_{F}$ could barely reach the lower limit of the experimental width range. While with $\alpha=2$, $\Gamma_{F}$ could reach the experimental observation when $g_{F\kappa\bar{\kappa}}$ is around $31$ GeV. Certainly with $\alpha=3$, the required $g_{F\kappa\bar{\kappa}}$ to reproduce the experimental width is further reduced to about $25$ GeV. For presenting the accurate quantities of the partial widths of different channels when the total width is fitted to experiments, we show the results in the third column in Table 2 with $g_{F\kappa\bar{\kappa}}=35$ GeV.
@@
From Table 1, we found that the U(3) symmetry expression of the $g^{\prime}_{\kappa K\eta}$ coupling consists of two destructive terms, which respectively correspond to the $\eta_{8}$ and $\eta_{1}$ parts in the mixing wave function. We adopt a relatively large value of $\theta=-10.7^{\circ}$ referring to the analysis in [62, 66]. Due to the comparable size of these two parts, the destructive interference leads to a small value of $g^{\prime}_{\kappa K\eta}$, which is sensitive to the mixing angle $\theta$. Hence, the partial width $\Gamma^{\prime}(F\to\eta\eta)$ calculated with $g^{\prime}_{\kappa K\eta}$ is the smallest, with its order of magnitude being one-thousandth of the partial widths of the other channels. Compared to it, $\Gamma(F\to\eta\eta)$ is larger by a factor of about $20$ due to the coupling $g_{\kappa K\eta}$ which is not suppressed by the destructive interference. But this coupling is still smaller than $g_{\kappa K\pi}$, which is contrary to the relation between $g_{K^{*}K\pi}$ and $g_{K^{*}K\eta}$: $g_{K^{*}K\pi^{\pm}}:g_{K^{*}K\eta}=1:(\sqrt{3}/\sqrt{2})\cos\theta$. Because the irreducible decomposition of $8\otimes 8$ include two different octet $8_{A}$ and $8_{S}$. Here $A$ and $S$ indicate the antisymmetric and symmetric features corresponding to the $K^{*}$ and $\kappa$ flavor wave functions respectively, which leads to different coupling relations. Therefore, the calculations treating the $f_{0}(1710)$ as a $K^{*}\bar{K}^{*}$ molecular state in Ref. [68] show that its decays into $\eta\eta$ and $\pi\pi$ channels are in the same order of magnitude. In contrast, in our case where $f_{0}(1370)$ is treated as a $\kappa\bar{\kappa}$ molecular state, it decays into these two channels with the same triangle diagram mechanism exchanging $K$, yet their partial widths differ by one to two orders of magnitude.
The $K\bar{K}\pi\pi$ channel is the only one proceeding through tree-level decay. However, the partial width of this channel is much smaller than that of the $K\bar{K}$, $\pi\pi$ and $4\pi$ channels, which arise from loop-level decays. This phenomenon might be mainly attributed to the very limited $K\bar{K}\pi\pi$ phase space around $\sqrt{s}=1.37$ GeV compared to the other channels, which is near the $\kappa\bar{\kappa}$ threshold. Consequently, around this energy region, the $K\bar{K}\pi\pi$ width is sensitive to the initial decay mass $\sqrt{s}$.
@@
Table 3: Decay widths of $f_{0}(1370)\to 4\pi$ with $g_{F\kappa\bar{\kappa}}=35$ GeV. Exclusive widths for each diagram and different intermediate channels are also presented. The notation [ex, $R\to ff$] in the second column specifies the exchange particle “ex” between the $\kappa$ pair and the intermediate resonance “RR” decaying to final states “ffff”. The decay widths are given in three different choices of the cutoff parameters $\alpha=1/2/3$. Decay widths are evaluated at initial energy $\sqrt{s}=1.37$ GeV and given in unit of MeV.
@@
With $g_{F\kappa\bar{\kappa}}=35$ GeV, the partial widths of different specific $K\bar{K}\pi\pi$ final states at $\sqrt{s}=1.37$ GeV are as follows: $\Gamma(K^{+}K^{-}\pi^{+}\pi^{-})=\Gamma(K^{0}\bar{K}^{0}\pi^{+}\pi^{-})=108$ keV; $\Gamma(K^{+}K^{-}\pi^{0}\pi^{0})=\Gamma(K^{0}\bar{K}^{0}\pi^{0}\pi^{0})=54$ keV; $\Gamma(K^{+}\bar{K}^{0}\pi^{0}\pi^{-})=\Gamma(K^{0}K^{-}\pi^{+}\pi^{0})=0.53$ keV. For the states with two charged $\pi$, the charged couplings $g_{\kappa K\pi^{\pm}}$ is larger than the neutral couplings by a factor of $\sqrt{2}$, while for the states including $\pi^{0}\pi^{0}$, there are another exchanged diagram and an extra symmetry factor $1/2$, finally leading to such a relation. For the states including only one charged pion, there are two diagrams respectively from the two different $\kappa\bar{\kappa}$ components which interfere destructively due to the $SU(3)$ relations listed in Table 1, resulting in a very small partial width for these states. Such a pattern is different from the direct decay to $K\bar{K}\pi\pi$, like the data in $J/\psi$ decay [8]. The same decay pattern among these three kinds of $K\bar{K}\pi\pi$ states should also appear in the $K^{*}\bar{K}^{*}\to(K\pi)(\bar{K}\pi)$ process due to the same flavor wavefunctions of $K^{*}$ and $\kappa$.
@@
It is worthy to mention that in our framework, the decay widths to $K\bar{K}$, $\pi\pi$, $\eta\eta$ and $K\bar{K}\pi\pi$ are sensitive to the coupling $g_{\kappa K\pi}$ because they are essentially proportional to the fourth power of this coupling. In fact, the $|\tilde{g}_{\kappa K\pi}|$ is determined to be $4.4$ [69] and $3.81$ [61] in two different works respectively, while such a small difference in this coupling can lead to a variation in these partial widths by nearly a factor of two.
@@
The partial widths of different specific $4\pi$ final states at $\sqrt{s}=1.37$ GeV are shown in Table 3, where we also present the exclusive contributions from different diagrams and intermediate states. The ratios between different $4\pi$ states are closely related to the decay mechanisms, as the intermediate states $\rho\rho$ and $\sigma\sigma$ have distinct features in this decay process. These differences can reflect the decay dynamics of the $f_{0}(1370)$ and thereby reveal its structure. Review the respective amplitudes for $2\pi^{+}2\pi^{-}$ and $\pi^{+}\pi^{-}2\pi^{0}$ as well as the results at $\sqrt{s}=1.37$ GeV presented in Table 3, we found that in our framework, the contributions from the $\sigma\sigma\to 2\pi^{+}2\pi^{-}$ are much larger than those from $\rho^{0}\rho^{0}\to 2\pi^{+}2\pi^{-}$. The widths of the process $\sigma\sigma\to\pi^{+}\pi^{-}2\pi^{0}$ is smaller than that of $\sigma\sigma\to 2\pi^{+}2\pi^{-}$ even though the symmetry factor for $\pi^{+}\pi^{-}2\pi^{0}$ is larger. This is mainly because there is no identical particle exchange diagram in the $\sigma\sigma\to\pi^{+}\pi^{-}2\pi^{0}$ process where the two $\pi^{0}$ are from the same $\sigma$. In contrast, the process $\sigma\sigma\to 2\pi^{+}2\pi^{-}$ does include the contributions from such an exchange diagram. After the analysis it is found that the amplitudes for $\rho\rho$ decay to $2\pi^{+}2\pi^{-}$ and $\pi^{+}\pi^{-}2\pi^{0}$ are the same then their widths only differ by the symmetry factor, which bring the relation of $2\Gamma(\rho\rho\to 2\pi^{+}2\pi^{-})=\Gamma(\rho\rho\to\pi^{+}\pi^{-}2\pi^{0})$ as listed in Table 3.
@@
Hence, if the structure of $f_{0}(1370)$ is more likely to decay via the $\sigma\sigma$ intermediate state, the width of the $2\pi^{+}2\pi^{-}$ channel will be approximately twice that of the $\pi^{+}\pi^{-}2\pi^{0}$ channel. In contrast, if the $\rho\rho$ intermediate state dominates, the width of the $\pi^{+}\pi^{-}2\pi^{0}$ will be almost twice that of the $2\pi^{+}2\pi^{-}$ channel. The inclusions of both types of processes will cause the ratio of $\Gamma_{2\pi^{+}2\pi^{-}}/\Gamma_{\pi^{+}\pi^{-}2\pi^{0}}$ to fall between these two limits. Comparing this ratio to experimental measurements can help clarify the decay mechanism of $f_{0}(1370)$, examining our scenario. Similarly, the ratio between the widths of the $4\pi^{0}$ state and the other $4\pi$ states can be used to estimate the contributions from the $\rho\rho$ intermediate state, as only the $\sigma\sigma$ intermediate state can contribute to the $4\pi^{0}$ final state.
@@
The decay pattern for a broad state is not as accurate to define at a fixed energy as the narrow state, owing to its wide span mass distribution. The partial widths of some channels might be sensitive to the initial mass $\sqrt{s}$ over its distribution where some relevant thresholds are located in. Hence, the well-defined branching ratios used to compare different decay channels for a broad state, are generally obtained by integrating over its mass distribution, a process typically done in the experimental analysis. For a further reliable study on the decay pattern of the $f_{0}(1370)$, we present the dependence of the partial widths of different channels on the initial energy $\sqrt{s}$. We plot the partial widths of different channels varying along with the initial energy $\sqrt{s}$ with fixed initial coupling $g_{F\kappa\bar{\kappa}}=35$ GeV and cut-off parameter $\alpha=2$ in Fig. 3.

### v2

@@ II Formalism
The constant $1/2$ is the symmetry factor on account of the two identical particles $\pi^{0}$.
For the $\sigma\sigma\to 4\pi^{0}$ process, the amplitudes $\mathcal{M}^{S}_{\sigma}$ and $\mathcal{M}^{V}_{\sigma}$ for $2\pi^{+}2\pi^{-}$ could be directly adopted but need to further considering all possible exchanges of the $\pi^{0}$. The decay width of the $4\pi^{0}$ state is
@@
The constant $1/4!$ is the symmetry factor on account of the four identical particles $\pi^{0}$.
For the $K\bar{K}\pi\pi$ channel from the tree-level decay $f_{0}(1370)\to\kappa\bar{\kappa}\to(K\pi)(\bar{K}\pi)$ shown in Fig. 1(d), there are six different final states specifically. The $K^{+}K^{-}\pi^{+}\pi^{-}$ and $K^{0}\bar{K}^{0}\pi^{0}\pi^{0}$ can only originate from the $\kappa^{0}\bar{\kappa}^{0}$ component; the $K^{0}\bar{K}^{0}\pi^{+}\pi^{-}$ and $K^{+}K^{-}\pi^{0}\pi^{0}$ can only originate from the $\kappa^{+}\kappa^{-}$ component; while the $K^{0}K^{-}\pi^{+}\pi^{0}$ and $K^{+}\bar{K}^{0}\pi^{0}\pi^{-}$ can originate from both the $\kappa^{0}\bar{\kappa}^{0}$ and $\kappa^{+}\kappa^{-}$ components.
@@
where $s_{ab}$ and $s_{cd}$ are respectively the invariant mass squares of the $K\pi$ pairs coming from $\kappa$ and $\bar{\kappa}$. The amplitudes of other final states can be obtained by replacing the couplings in $\mathcal{M}_{K^{+}K^{-}\pi^{+}\pi^{-}}$ and carefully including all the possible diagrams. The combinations of $K\pi$ are definite for all the $K\bar{K}\pi\pi$ channels except for $K^{+}K^{-}\pi^{0}\pi^{0}$ and $K^{0}\bar{K}^{0}\pi^{0}\pi^{0}$ due to the two identical particles $\pi^{0}$. Hence, for these two final states, we need to include another diagram obtained by exchanging these two $\pi^{0}$ and when calculating the decay width, the symmetry factor $1/2$ should be considered on account of the double counting in the phase space.
@@ III Numerical results and discussions
Firstly, we present the results calculated with the coupling $g_{F\kappa\bar{\kappa}}$ obtained by Weinberg criterion. The partial decay widths of different channels including $K\bar{K}$, $\pi\pi$, $\eta\eta$, $4\pi$ and $K\bar{K}\pi\pi$ are shown in Table 2. These are the primary strong decay channels of $f_{0}(1370)$, which are approximately summed up to the total width of $f_{0}(1370)$: the $\Gamma_{F}$ in Table 2. We present results for two different couplings $g_{\kappa K\eta}$ and $g^{\prime}_{\kappa K\eta}$. For clarity, we denote the result associated with $g^{\prime}_{\kappa K\eta}$ as $\Gamma^{\prime}$. We approximate the total $4\pi$ width by summing the $\sigma\sigma$ and $\rho\rho$ contributions, under the assumption that other intermediate states are negligible. In our framework, the $f_{0}(1370)$ decaying to the $K\bar{K},\pi\pi,\eta\eta$ and $4\pi$ channels are via the triangle loop diagrams, which are dependent on the cut-off parameters $\alpha$. Hence, we also present the results with different $\alpha$ in Table 2, to study the cut-off dependence. From Table 2, we can see that when $\alpha$ varies from $2$ to $3$, the decay widths are relatively stable. These decay channels have been observed in experiments except the $K\bar{K}\pi\pi$ channel, which lacks measurements up to now.
Regardless of whether the components are broad or narrow, they are nearly free particles inside the molecular state in the weak-binding limit. Decay through the almost free components implies that the width of the bound state is expected to be of the same order of magnitude as the width of its components. This argument is consistent with the case of $T^{+}_{cc}$, whose width is almost saturated by the decay via $D^{*}D\to DD\pi$ and is analyzed to be about $48$ keV [67], and the width of $D^{*}$ is $(83.4\pm 1.8)$ keV [8]. Hence, aligning with this argument the broad state $f_{0}(1370)$ is reasonably suspected to be a molecular state composed of $\kappa\bar{\kappa}$. The pole width $\Gamma=-2\,\mathrm{Im}\,\sqrt{s_{\text{pole}}}$ for $f_{0}(1370)$ is estimated to be around $500$ MeV according to its pole position, and its Breit-Wigner width is measured to be between $200$ MeV and $500$ MeV depending on the models and the reactions. However, as shown in Table 2, with the $g_{F\kappa\bar{\kappa}}=13.4$ GeV determined by the Weinberg criterion and the cut off parameter $\alpha=2$, the sum of all decay channels results in a total width $\Gamma_{F}=39.6/\Gamma^{\prime}_{F}=37.7$ MeV, which is much smaller than all the measured widths of $f_{0}(1370)$ collected in the RPP [8]. Recall that the Weinberg criterion is derived from the case of weakly bound state consisting of two stable or nearly stable components. In contrast, the component $\kappa$ in our case is broad as about $400$ MeV, which implies that the decay coupling $g_{F\kappa\bar{\kappa}}$ might still be underestimated even though a complex mass are used in the Weinberg criterion.
@@
We realize that it is complex to figure out the coupling between $f_{0}(1370)$ and its broad components, let alone this molecular state might essentially be composed of four stable particles $K\bar{K}\pi\pi$. Therefore, we take this couplings as a parameter to fit the total width of $f_{0}(1370)$ measured in experiments. In Fig. 2, we show the total width of $f_{0}(1370)$ at $\sqrt{s}=1.37$ GeV with $g_{F\kappa\bar{\kappa}}$ as a parameter. We also present the results with different cut-off parameters $\alpha$ from $1\sim 3$. Notice that $g_{F\kappa\bar{\kappa}}$ is a common coupling for all the decay channels, its change will not affect the ratios among the decay channels. In other words, the decay pattern is independent of this coupling. The light purple band in Fig. 2 indicates the range of the Breit-Wigner width of $f_{0}(1370)$ measured in experiments. The vertical red dashed line indicates the value of $g_{F\kappa\bar{\kappa}}$ from the Weinberg criterion using a complex threshold. We can see that when cut-off parameter $\alpha=1$, even the coupling $g_{F\kappa\bar{\kappa}}$ is taken as large as $45$ GeV, the total width $\Gamma_{F}$ could barely reach the lower limit of the experimental width range. While with $\alpha=2$, $\Gamma_{F}$ can reach the experimental observation when $g_{F\kappa\bar{\kappa}}$ is around $31$ GeV. Certainly with $\alpha=3$, the required $g_{F\kappa\bar{\kappa}}$ to reproduce the experimental width is further reduced to about $25$ GeV. For presenting the accurate quantities of the partial widths of different channels when the total width is fitted to experiments, we show the results in the third column in Table 2 with $g_{F\kappa\bar{\kappa}}=35$ GeV.
@@
From Table 1, we find that the U(3) symmetry expression of the $g^{\prime}_{\kappa K\eta}$ coupling consists of two destructive terms, which respectively correspond to the $\eta_{8}$ and $\eta_{1}$ parts in the mixing wave function. We adopt a relatively large value of $\theta=-10.7^{\circ}$ referring to the analysis in [62, 66]. Due to the comparable size of these two parts, the destructive interference leads to a small value of $g^{\prime}_{\kappa K\eta}$, which is sensitive to the mixing angle $\theta$. Hence, the partial width $\Gamma^{\prime}(F\to\eta\eta)$ calculated with $g^{\prime}_{\kappa K\eta}$ is the smallest, with its order of magnitude being one-thousandth of the partial widths of the other channels. Compared to it, $\Gamma(F\to\eta\eta)$ is larger by a factor of about $20$ because the coupling $g_{\kappa K\eta}$ is not suppressed by the destructive interference. But $g_{\kappa K\eta}$ is still smaller than $g_{\kappa K\pi}$, which is contrary to the relation between $g_{K^{*}K\pi}$ and $g_{K^{*}K\eta}$: $g_{K^{*}K\pi^{\pm}}:g_{K^{*}K\eta}=1:(\sqrt{3}/\sqrt{2})\cos\theta$. Because the irreducible decomposition of $8\otimes 8$ include two different octet $8_{A}$ and $8_{S}$. Here $A$ and $S$ indicate the antisymmetric and symmetric features corresponding to the $K^{*}$ and $\kappa$ flavor wave functions respectively, which leads to different coupling relations. Therefore, the calculations treating the $f_{0}(1710)$ as a $K^{*}\bar{K}^{*}$ molecular state in Ref. [68] show that its decays into $\eta\eta$ and $\pi\pi$ channels are in the same order of magnitude. While with the same triangle diagram mechanism exchanging $K$, the partial widths of $f_{0}(1370)$ treated as a $\kappa\bar{\kappa}$ molecule to these two channels differ by one to two orders of magnitude.
The $K\bar{K}\pi\pi$ channel is the only one proceeding through tree-level decay. However, the partial width of this channel is much smaller than that of the $K\bar{K}$, $\pi\pi$ and $4\pi$ channels, which arise from loop-level decays. This phenomenon might be mainly attributed to the very limited $K\bar{K}\pi\pi$ phase space compared to the other channels around $\sqrt{s}=1.37$ GeV, which is near the $\kappa\bar{\kappa}$ threshold. Consequently, around this energy region, the $K\bar{K}\pi\pi$ width is sensitive to the center-of-mass energy $\sqrt{s}$.
@@
Table 3: Decay widths of $f_{0}(1370)\to 4\pi$ with $g_{F\kappa\bar{\kappa}}=35$ GeV. Exclusive widths for each diagram and different intermediate channels are also presented. The notation [ex, $R\to ff$] in the second column specifies the exchange particle “ex” between the $\kappa$ pair and the intermediate resonance “RR” decaying to final states “ffff”. The decay widths are given in three different choices of the cutoff parameters $\alpha=1/2/3$. Decay widths are evaluated at center-of-mass energy $\sqrt{s}=1.37$ GeV and given in unit of MeV.
@@
With $g_{F\kappa\bar{\kappa}}=35$ GeV, the partial widths of different specific $K\bar{K}\pi\pi$ final states at $\sqrt{s}=1.37$ GeV are as follows: $\Gamma(K^{+}K^{-}\pi^{+}\pi^{-})=\Gamma(K^{0}\bar{K}^{0}\pi^{+}\pi^{-})=108$ keV; $\Gamma(K^{+}K^{-}\pi^{0}\pi^{0})=\Gamma(K^{0}\bar{K}^{0}\pi^{0}\pi^{0})=54$ keV; $\Gamma(K^{+}\bar{K}^{0}\pi^{0}\pi^{-})=\Gamma(K^{0}K^{-}\pi^{+}\pi^{0})=0.53$ keV. For the states with two charged $\pi$, the charged couplings $g_{\kappa K\pi^{\pm}}$ is larger than the neutral couplings by a factor of $\sqrt{2}$. While for the states including $\pi^{0}\pi^{0}$, there are another exchanged diagram and an extra symmetry factor $1/2$, finally leading to such a relation. For the states including only one charged pion, there are two diagrams respectively from the two different $\kappa\bar{\kappa}$ components which interfere destructively due to the $SU(3)$ relations listed in Table 1, resulting in a very small partial width for these states. Such a pattern is different from the direct decay to $K\bar{K}\pi\pi$, such as the data in $J/\psi$ decay [8]. The same decay pattern among these three kinds of $K\bar{K}\pi\pi$ states should also appear in the $K^{*}\bar{K}^{*}\to(K\pi)(\bar{K}\pi)$ process due to the same relations of couplings in $K^{*}K\pi$ and $\kappa K\pi$.
@@
It is worth mentioning that in our framework, the decay widths to $K\bar{K}$, $\pi\pi$, $\eta\eta$ and $K\bar{K}\pi\pi$ are sensitive to the coupling $g_{\kappa K\pi}$ because they are essentially proportional to the fourth power of this coupling. In fact, the $|\tilde{g}_{\kappa K\pi}|$ is determined to be $4.4$ [69] and $3.81$ [61] in two different works respectively, while such a small difference in this coupling can lead to a variation in these partial widths by nearly a factor of two.
@@
The partial widths of different specific $4\pi$ final states at $\sqrt{s}=1.37$ GeV are shown in Table 3, where we also present the exclusive contributions from different diagrams and intermediate states. The ratios between different $4\pi$ states are closely related to the decay mechanisms, as the intermediate states $\rho\rho$ and $\sigma\sigma$ have distinct features in $4\pi$ decay. These differences can reflect the decay dynamics of the $f_{0}(1370)$ and thereby reveal its structure. Reviewing the respective amplitudes for $2\pi^{+}2\pi^{-}$ and $\pi^{+}\pi^{-}2\pi^{0}$ as well as the results at $\sqrt{s}=1.37$ GeV presented in Table 3, we find that in our framework, the contributions from the $\sigma\sigma\to 2\pi^{+}2\pi^{-}$ are much larger than those from $\rho^{0}\rho^{0}\to 2\pi^{+}2\pi^{-}$. The widths of the process $\sigma\sigma\to\pi^{+}\pi^{-}2\pi^{0}$ is smaller than that of $\sigma\sigma\to 2\pi^{+}2\pi^{-}$ even though the symmetry factor for $\pi^{+}\pi^{-}2\pi^{0}$ is larger. This is mainly because there is no identical particle exchange diagram in the $\sigma\sigma\to\pi^{+}\pi^{-}2\pi^{0}$ process where the two $\pi^{0}$ are from the same $\sigma$. In contrast, the process $\sigma\sigma\to 2\pi^{+}2\pi^{-}$ does include the contributions from such an exchange diagram. After the analysis it is found that the amplitudes for $\rho\rho$ decay to $2\pi^{+}2\pi^{-}$ and $\pi^{+}\pi^{-}2\pi^{0}$ are the same then their widths only differ by the symmetry factor, which bring the relation of $2\Gamma(\rho\rho\to 2\pi^{+}2\pi^{-})=\Gamma(\rho\rho\to\pi^{+}\pi^{-}2\pi^{0})$ as listed in Table 3.
@@
Hence, if the structure of $f_{0}(1370)$ is more likely to decay via the $\sigma\sigma$ intermediate state, the width of the $2\pi^{+}2\pi^{-}$ channel will be approximately twice that of the $\pi^{+}\pi^{-}2\pi^{0}$ channel. In contrast, if the $\rho\rho$ intermediate state dominates, the width of the $\pi^{+}\pi^{-}2\pi^{0}$ will be almost twice that of the $2\pi^{+}2\pi^{-}$ channel. Including both types of processes, the ratio of $\Gamma_{2\pi^{+}2\pi^{-}}/\Gamma_{\pi^{+}\pi^{-}2\pi^{0}}$ falls between these two limits. Comparing this ratio to experimental measurements can help clarify the decay mechanism of $f_{0}(1370)$, examining our scenario. Similarly, the ratio between the widths of the $4\pi^{0}$ state and the other $4\pi$ states can be used to estimate the contributions from the $\rho\rho$ intermediate state, as only the $\sigma\sigma$ intermediate state can contribute to the $4\pi^{0}$ final state.
@@
The decay pattern for a broad state is not as accurate to define at a fixed energy as the narrow state, owing to its wide span mass distribution. The partial widths of some channels might be sensitive to the center-of-mass energy $\sqrt{s}$ over its distribution where some relevant thresholds are located in. Hence, the well-defined branching ratios of different decay channels for a broad state, are generally obtained by integrating over its mass distribution, a process typically done in the experimental analysis. For a further reliable study on the decay pattern of the $f_{0}(1370)$, we present the dependence of the partial widths of different channels on the center-of-mass energy $\sqrt{s}$. We plot the partial widths of different channels varying along with the energy $\sqrt{s}$ with fixed coupling $g_{F\kappa\bar{\kappa}}=35$ GeV and cut-off parameter $\alpha=2$ in Fig. 3.